# Quantum Down Sampling Filter for Variational Auto-encoder

### v1

@@ 3.1.2 Windowing and Pooling for Resolution Reduction
Pooling techniques such as max pooling or average pooling play a critical role in reducing the spatial resolution of the input data while retaining important features. In max pooling, for example, the maximum value in a small region (or window) of the input image is selected, effectively reducing the image’s resolution while retaining the most significant features within each region. Average pooling operates similarly but instead computes the average value of each window. These pooling operations are vital for reducing computational complexity while preserving essential information for later stages of the model. Adaptive pooling is another technique often used in encoders to ensure that feature maps have consistent sizes, regardless of the input size. Adaptive pooling adjusts the pooling operation to produce fixed-sized output maps, which is particularly useful when working with images of varying resolutions. This resolution reduction helps the model to focus on more abstract features without losing key information, thus enabling more efficient learning in the latent space.
@@ 3.2 Latent Space Encoding
The proposed architecture effectively combines quantum and classical components, leveraging quantum-enhanced feature extraction for downstream Artificial Neural Network (ANN) processing.
@@ 3.3 Decoder
The decoder in a Variational Auto-encoder (VAE) is designed to transform the lower-resolution latent variable $z$, produced by the encoder, back into a high-resolution data space, reconstructing the input data. The decoder typically consists of multiple layers, including fully connected and transposed convolution layers, which progressively reconstruct the input data’s spatial and feature dimensions. The core function of the decoder is to ensure that the reconstructed data closely matches the original input, enabling the model to learn both the global structure and local details of the data. As with the encoder, the decoder is trained to minimize the reconstruction error, often using a loss function that combines both the likelihood of the data given the latent variables and a regularization term that encourages smoothness in the latent space representation.
@@ 3.3.1 Latent Variable Transformation
The decoding process begins by passing the latent variable $z$ through a fully connected layer to map it into a higher-resolution feature space. This transformation creates a tensor that matches the expected dimensions for the next layers in the decoder, such as the transposed convolution layers. The fully connected layer enables the decoder to effectively learn a mapping from the compact latent space back to the data space, ensuring that relevant information is maintained during this transformation.
@@ 3.3.2 Transposed Convolutions
Once the latent variable is transformed into the higher-resolution space, it is passed through a series of transposed convolution (also known as de-convolution) layers. These layers up-sample the tensor, progressively restoring the original spatial dimensions of the input data. Each transposed convolution layer learns to refine the spatial resolution of the image while preserving important features. Transposed convolution layers have been shown to be highly effective in generating high-resolution outputs from low-resolution latent representations. They are capable of learning hierarchical features of the data, capturing both local details and global structure during the reconstruction process. By applying nonlinear activation functions, such as ReLU, between the transposed convolution layers, the model introduces the necessary complexity to capture intricate patterns in the data.
@@ 3.4 Quantum Computing in Variational Auto-encoder
To address these limitations, we propose the integration of quantum computing into the VAE encoder. Quantum computing has shown promise in enhancing generative models due to its ability to represent information in quantum states. Unlike classical models, which rely on deterministic mappings, quantum encoders can utilize superposition and entanglement to encode data into a potentially much higher-resolution space, enabling the model to learn richer, more compact latent representations. Quantum circuits can efficiently explore large state spaces, leveraging the Hilbert space to represent data points in ways that are not possible with classical methods [14]. Mathematically, quantum encoders can represent the latent variables $\mathbf{z}$ as quantum states $|\psi\rangle\in\mathcal{H}$, where $\mathcal{H}$ is the Hilbert space. A quantum encoder might take the form of a unitary operator $U_{\theta}$ applied to an initial quantum state $|\mathbf{x}\rangle$, mapping the input data $\mathbf{x}$ to a quantum state that encodes the latent representation:
@@
where $U_{\theta}$ is a quantum operation that performs the encoding. This encoding could be leveraged to provide a quantum-enhanced latent space that could more efficiently represent the complex structures within the data. Quantum states encoded in this way can capture correlations in data that are hard to model classically, such as in the case of high-resolution images. The advantage of quantum encoding lies in its ability to utilize the inherent properties of quantum mechanics, such as quantum entanglement, which allows for more efficient representation of correlations within data. Additionally, quantum superposition enables the parallel encoding of multiple latent variables simultaneously, potentially leading to a richer and more compact latent space. This could potentially address one of the critical issues in high-resolution data reconstruction, where classical VAEs often struggle with blurred images due to the loss of important data features in the latent space. Moreover, quantum-based encoders have potential to improve the feature extraction process. This could enable the model to perform better in tasks such as image reconstruction, de-noising, and anomaly detection.
@@ 4 Proposed Methodology
In this work, we introduce a novel down sampling filter for Quantum Variational Auto-encoders (Q-VAE) that relies exclusively on quantum encoding only, for feature extraction. The paper investigates three distinct architectures for Variational Auto-encoders (VAEs), each employing different encoding mechanisms to evaluate their feature extraction capabilities: Quantum VAE (Q-VAE), Classical VAE (C-VAE), and Direct Passing VAE (CDP-VAE). The purpose of this comparison is to evaluate the effectiveness of quantum computing, classical neural networks, and simplified encoding strategies the capture and processes data for downstream reconstruction. While each model shares a common decoder architecture, the encoders differ in their approach, enabling the study of how each influences the encoding process and the quality of the reconstructed output.
@@ 4.1 Quantum VAE (Q-VAE) Encoder
The Quantum VAE (Q-VAE) integrates quantum computing into the encoding process, replacing the traditional classical neural network encoder with a quantum circuit as shown in Figure 1. The quantum encoder maps the input data into a quantum state representation, leveraging quantum superposition to capture complex relationships in the data. By encoding the input into quantum states, Q-VAE is able to extract features from data that may be difficult for classical methods to model effectively, especially in high-resolution or non-linear data distributions. The quantum-enhanced feature extraction is paired with a classical CNN decoder that reconstructs the input data, combining the strengths of quantum processing for feature extraction with classical techniques for data reconstruction.
Figure 1: Architecture for Variational Quantum Auto-encoder.
The encoder in our model compresses input data by encoding it into a lower-resolution latent space representation. It consists of a quantum circuit as shown in Figure 2, integrated with a classical fully connected layer.
Mathematically, the encoder transforms the input data $X$ into two key outputs: the mean $\mu$ and the log variance $\log(\sigma^{2})$ of the latent space distribution:
$z\sim\mathcal{N}(\mu,\sigma^{2}),\quad\mu=f_{\text{encoder}}(X),\quad\log(\sigma^{2})=g_{\text{encoder}}(X),$ (5)
where $f_{\text{encoder}}$ and $g_{\text{encoder}}$ are parameterized by the quantum-classical network.
@@
The quantum encoding network, illustrated in Figure 2, uses a quantum circuit designed to process 32x32 pixel images (1024 pixels in total). This circuit operates on 4 qubits, encoding 4 pixels at a time. For a window of 4 pixels $[x_{1},x_{2},x_{3},x_{4}]$, the first pixel ($x_{1}$) is encoded into the quantum state using a rotation gate:
@@
The measurement on qubit 0 yields the compressed representation of the input. This procedure reduces the resolutionity of the input image from 1024 pixels to 256 measured values as shown in figure 2. After the quantum circuit processes the input image and extracts quantum features, the output of the quantum circuit, as shown in figure 3, is passed to fully connected layer. This transition enables the integration of quantum-encoded features with classical deep learning architectures. The quantum circuit produces an output, typically in the form of a reduced-resolution feature map or a vector. Let the output from the quantum circuit be denoted as $q_{\text{out}}$, where:
@@
4.2 Classical VAE (C-VAE) Encoder
The Classical VAE (C-VAE) uses a traditional fully connected neural network as the encoder, which processes the input data through a series of hidden layers to output the mean ($\mu$) and log variance ($\log(\sigma^{2})$) of the latent space distribution. This distribution is then sampled to obtain the latent variable $z$, which represents the encoded data in a lower-resolution space. While the C-VAE relies entirely on classical neural networks for encoding, it remains highly effective in modeling data distributions and is widely used in practical applications. The decoder, typically a CNN, reconstructs the input from this low-resolution latent space, generating an output image or data structure that approximates the original input.
4.3 Classical Direct Passing VAE (CDP) Architecture
Figure 4: Direct Passing CDP Variational Auto encoder Architecture
Mathematically, the CDP-VAE encoder operates as follows:
Figure 5: USPS image Direct Passing filter conversion to 16x16
$[x_{\text{reduced}}=\text{Resample}(x),$ (13)
@@ 4.4 Shared Decoder Architecture
Despite the differences in their encoding mechanisms, all three models, Q-VAE, C-VAE, and CDP-VAE, share a common decoder architecture. This decoder is responsible for reconstructing the input data from the latent space representation. The decoder begins by transforming the latent variable \( z \) into a higher-resolution space through fully connected layers. Then, a series of transposed convolution (de-convolution) layers are used to progressively up-sample the data, restoring the original spatial dimensions. Finally, the output is passed through a sigmoid activation function to ensure that the pixel values are within the valid range (e.g,, between 0 and 1 for image data). This shared decoder allows for a fair comparison of the encoding techniques, with the focus on how the different encoders influence the quality of the reconstructed image. .
@@ 5 Numerical Simulation
5.1 MNIST Dataset PreProcessing
@@
    Resizing to $16\times 16$: After cropping, the image is resized to $16\times 16$ pixels. This step reduces the image resolution, making it easier for the model to process while still retaining key features of the digit. The resizing step uses an interpolation technique that preserves the overall structure of the digit. This process will help us to get same resolution dataset like USPS to see if MNIST also show improvement or not.
@@
    Resizing to $32\times 32$: Finally, the image is resized to $32\times 32$ pixels. This larger size is used for feeding the images into the neural network model, which is optimized to process higher resolution images. This resizing allows the model to capture finer details of the digits, enhancing its ability to reconstruct and classify the images effectively[15].
@@
5.2 USPS Dataset PreProcessing
@@
The interpolation method used during resizing adjusts the pixel values to fit the larger image grid, enabling the model to be tested with standard image input sizes, such as 32x32.
@@ 5.3 Experimental Setup
In this section, we describe the hyper-parameters and settings used for training the models. These hyper-parameters were carefully selected to ensure stable training while optimizing performance for both models.
@@ 5.3.1 Hyper-parameters
In our training process, several key hyper-parameters were used to ensure efficient learning and model convergence. The learning rate was set to $0.001$ for all the models. This learning rate controls how much the model’s weights are updated in response to the computed gradients, providing a stable training process while allowing sufficient progress over the course of the training. The models were trained for 200 epochs, with each epoch representing one full pass of the data set through the model. This duration was chosen to ensure that the models had enough time to learn the underlying features of the data, while regular monitoring of the validation loss helped prevent over-fitting. A batch size of $400$ was used, which means that $400$ images were processed together before updating the model parameters. This batch size was selected to balance computational efficiency and the stability of gradient estimation, helping smooth gradient updates and ensuring more stable training. For optimization, we used the Adam optimizer, which is known for its adaptive learning rate that adjusts based on the first and second moments of the gradients. The Adam optimizer is particularly effective for models with many parameters and large datasets. The loss function used was the Binary Cross-Entropy Loss (BCELoss), with the argument reduction = ’sum’, which ensures that the individual losses across all pixels are summed rather than averaged. This choice emphasizes the total error in the reconstruction of the images. To visualize the progress of the model over time, the variables q_grid_images, cdp_grid_images and c_grid_images were used to store a grid of reconstructed images at various epochs. These images allowed us to visually assess how well the models were reconstructing the images as training progressed. These configurations were chosen to ensure effective model training, allowing the models to learn from the data efficiently and perform well in both reconstruction and evaluation tasks.
6 Performance Evaluation and Discussion
To evaluate the performance of the proposed models, we used a combination of several key metrics: the Fréchet Inception Distance (FID) score[17], training and testing loss, and Mean Squared Error (MSE)[18] The results are shown in table 1 below:
@@ 6 Performance Evaluation and Discussion
Each of these metrics provides distinct insights into the model’s ability to generate high-quality images and generalize to new data: Fréchet Inception Distance (FID) Score measures the similarity between the distribution of real and generated images. A lower FID score indicates that the generated images are closer in distribution to the real images, reflecting higher image fidelity and quality. The FID score is particularly useful for assessing the realism of generated images in generative models. Mean Squared Error (MSE) measures the average squared difference between the generated and real images. A lower MSE indicates that the generated images are closer to the real images in terms of pixel-wise accuracy. This metric is commonly used to evaluate the reconstruction quality in generative models, with lower values indicating better performance. By combining these metrics, we gain a comprehensive understanding of how well our models are performing in terms of both image quality and generalization ability.
The FID score was computed for 1) images reconstructed from the input images and 2) images generated from random noise, giving us insight into how well the models perform in generating realistic images from these two different sources of input. The performance of the C-VAE, Q-VAE, and CDP-VAE on the MNIST dataset was assessed by calculating the FID score using latent space representations as shown in Figure 8:
@@
The C-VAE achieves an FID score of 40.7, indicating a reasonable ability to reconstruct images. The Q-VAE, on the other hand, leverages quantum gates such as RY rotations and Pauli-Z measurements to encode the data into a quantum state, achieving a superior FID score of 37.3. This improved score suggests that the quantum model produces more faithful image reconstructions, capturing complex relationships in the data more effectively than the classical model. Finally, the CDP-VAE, demonstrates a slight improvement over the C-VAE with an FID score of 39.7. Although this approach is less complex, it still lacks the advanced feature extraction capabilities of the Q-VAE, which ultimately leads to lower reconstruction fidelity. For the USPS dataset, the Q-VAE achieved an impressive FID score of 57.6, indicating high-quality image reconstructions and strong performance in terms of both fidelity and distribution similarity to the original dataset. This result highlights the effectiveness of the Q-VAE in handling simpler grayscale datasets, where it demonstrates superior feature extraction and reconstruction capabilities.
@@
The CDP-VAE is a VAE where, instead of using an encoder-decoder architecture with learned transformations, the input data is directly passed through with minimal transformation using windowing/ pooling. This means the model relies on a simpler mechanism where the input features, such as image pixels, are either passed as-is or undergo basic operations. Despite its simplicity, the CDP-VAE still achieved relatively good performance on the USPS dataset, with an FID score of 66.1 as shown in Figure 9.

### v2

@@ 3.1.2 Windowing and Pooling for Resolution Reduction
Pooling techniques such as max pooling or average pooling play a critical role in reducing the spatial resolution of the input data while retaining important features. In max pooling, for example, the maximum value in a small region (or window) of the input image is selected, effectively reducing the image’s resolution while retaining the most significant features within each region. Average pooling operates similarly, but instead computes the average value of each window. These pooling operations are vital for reducing computational complexity while preserving essential information for later stages of the model. Adaptive pooling is another technique often used in encoders to ensure that feature maps have consistent sizes, regardless of the input size. Adaptive pooling adjusts the pooling operation to produce fixed-sized output maps, which is particularly useful when working with images of varying resolutions. This resolution reduction helps the model focus on more abstract features without losing key information, enabling more efficient learning in the latent space.
@@ 3.2 Latent Space Encoding
The proposed architecture effectively combines quantum and classical components, leveraging quantum-enhanced feature extraction for downstream Neural Network processing.
@@ 3.3 Decoder
The decoder in a Variational Auto-encoder (VAE) is designed to transform the lower-resolution latent variable $z$, produced by the encoder, back into a high-resolution data space, reconstructing the input data. The decoder typically consists of multiple layers, including fully connected and transposed convolution layers, which progressively reconstruct the input data’s spatial and feature dimensions. The core function of the decoder is to ensure that the reconstructed data match closely the original input, enabling the model to learn both the global structure and the local details of the data. As with the encoder, the decoder is trained to minimize the reconstruction error, often using a loss function that combines both the likelihood of the data given the latent variables and a regularization term that encourages smoothness in the latent-space representation.
@@ 3.3.1 Latent Variable Transformation
The decoding process begins by passing the latent variable $z$ through a fully connected layer to map it to a higher resolution feature space. This transformation creates a tensor that matches the expected dimensions for the next layers in the decoder, such as the transposed convolution layers. The fully connected layer enables the decoder to effectively learn a mapping from the compact latent space back to the data space, ensuring that relevant information is maintained during this transformation.
@@ 3.3.2 Transposed Convolutions
Once the latent variable is transformed into the higher-resolution space, it is passed through a series of transposed convolution (also known as de-convolution) layers.These layers up-sample the tensor, progressively restoring the original spatial dimensions of the input data. Each transposed convolution layer learns to refine the spatial resolution of the image while preserving important features. Transposed convolution layers have been shown to be highly effective in generating high-resolution outputs from low-resolution latent representations. They are capable of learning hierarchical features of the data, capturing both local details and global structure during the reconstruction process. By applying non-linear activation functions, such as ReLU, between the transposed convolution layers, the model introduces the necessary complexity to capture intricate patterns in the data.
@@ 3.4 Quantum Computing in Variational Auto-encoder
Quantum computing has shown promise in enhancing generative models due to its ability to represent information in quantum states. Unlike classical models, which rely on deterministic mappings, quantum encoders can utilize superposition and entanglement to encode data into a potentially much larger computational space, enabling the model to learn richer and more compact latent representations [20]. Quantum circuits can efficiently explore large state spaces, leveraging the Hilbert space to represent data points in ways that are not possible with classical methods [15]. Mathematically, quantum encoders can represent the latent variables $\mathbf{z}$ as quantum states $|\psi\rangle\in\mathcal{H}$, where $\mathcal{H}$ is the Hilbert space. A quantum encoder might take the form of a unitary operator $U_{\theta}$ applied to an initial quantum state $|\mathbf{x}\rangle$, mapping the input data $\mathbf{x}$ to a quantum state that encodes the latent representation that map the medium-dimensional data into a more compact, or lower-dimensional space.
@@
where $U_{\theta}$ is a quantum operation that performs the encoding. This encoding could be leveraged to provide a quantum-enhanced latent space that could more efficiently represent complex structures within the data. Quantum states encoded in this way can capture correlations in data that are hard to model classically, such as in the case of medium-resolution images.
The advantage of quantum properties has potential to improve tasks like image reconstruction and anomaly detection, overcoming challenges like blurred images often encountered with classical VAEs.
@@ 4 Proposed Methodology
In this work, we introduce a novel down-sampling filter for quantum variational auto-encoders (Q-VAE) that relies exclusively on quantum encoding for feature extraction. The paper investigates three distinct architectures for Variational Auto-encoders (VAEs), each employing different encoding mechanisms to evaluate their feature extraction capabilities: Quantum VAE (Q-VAE), Classical VAE (C-VAE), and Direct Passing VAE (CDP-VAE). The framework of these models has been presented in Figure 1. The goal of this comparison is to assess the effectiveness of quantum computing, classical neural networks, and simplified encoding strategies in capturing and processing data for subsequent reconstruction. While each model shares a common decoder architecture, the encoders differ in their approach, enabling the study of how each influences the encoding process and the quality of the reconstructed output.
Figure 1: Flowchart for (a) C-VAE, (b) CDP-VAE and (c) Q-VAE.
4.1 Classical VAE (C-VAE) Encoder
The Classical VAE (C-VAE) uses a traditional fully connected neural network as the encoder, which processes the input data through a series of hidden layers to output the mean ($\mu$) and log variance ($\log(\sigma^{2})$) of the latent space distribution. This distribution is then sampled to obtain the latent variable $z$, which represents the encoded data in a lower-resolution space. While the C-VAE relies entirely on classical neural networks for encoding, it remains highly effective in modeling data distributions and is widely used in practical applications. The decoder, typically a CNN, reconstructs the input from this low-resolution latent space, generating an output image or data structure that approximates the original input, also shown in Figure 1.
4.2 Classical Direct Passing (CDP-VAE) Encoder
Figure 2: USPS image Direct Passing filter conversion to 16x16
$[x_{\text{reduced}}=\text{Resample}(x),$ (5)
@@ 4.3 Quantum VAE (Q-VAE) Encoder
The Quantum VAE (Q-VAE) integrates quantum computing into the encoding process, replacing the traditional classical neural network encoder with a quantum circuit as shown in Figure 1. The quantum encoder maps the input data into a quantum state representation, leveraging quantum superposition to capture complex relationships in the data. Quantum encoding enhances performance by converting the input into quantum states, enabling the Q-VAE to apply rotations to the data. The quantum-enhanced feature extraction is paired with a classical CNN decoder that reconstructs the input data, combining the strengths of quantum processing for feature extraction with classical techniques for data reconstruction.
The Q-VAE encoder compresses input data by encoding it into a lower-resolution latent space representation. It consists of a quantum circuit as shown in Figure 3, integrated with a classical fully connected layer.
The encoder transforms the input data $X$ into two key outputs: the mean $\mu$ and the log variance $\log(\sigma^{2})$ of the latent space distribution:
@@
The quantum encoding network, illustrated in Figure 2, uses a quantum encoding designed to process 16x16 pixel images that is the output of windowing pooling layer.
@@
After the quantum circuit processes the input image and extracts quantum features, the output of the quantum circuit, as shown in figure 4, is passed to fully connected layer. This transition enables the integration of quantum-encoded features with classical deep learning architectures. The quantum circuit produces an output, typically in the form of a reduced-resolution feature map or a vector. Let the output from the quantum circuit be denoted as $q_{\text{out}}$, where:
@@ 4.4 Shared Decoder Architecture
Despite the differences in their encoding mechanisms, all three models, C-VAE, CDP-VAE and Q-VAE share a common decoder architecture. This decoder is responsible for reconstructing the input data from the latent space representation. The decoder begins by transforming the latent variable \( z \) into a medium-resolution space through fully connected layers. Then, a series of transposed convolution (de-convolution) layers are used to progressively up-sample the data, restoring the original spatial dimensions. Finally, the output is passed through a sigmoid activation function to ensure that the pixel values are within the valid range (e.g,, between 0 and 1 for image data). This shared decoder allows for a fair comparison of the encoding techniques, with the focus on how the different encoders influence the quality of the reconstructed image.
@@ 5 Numerical Simulation
5.1 MNIST Dataset Pre-Processing
@@
    Resizing to $16\times 16$: After cropping, the image is resized to $16\times 16$ pixels. This step reduces the image resolution, making it easier for the model to process while still retaining key features of the digit. This process will help us to get same resolution dataset like USPS to see if MNIST also show improvement or not.
@@
    Resizing to $32\times 32$: Finally, the image is resized to $32\times 32$ pixels. This larger size is used for feeding the images into the neural network model, which is optimized to process medium resolution images. This resizing allows the model to capture finer details of the digits, enhancing its ability to reconstruct and classify the images effectively[16].
@@
5.2 USPS Dataset Pre-Processing
@@
USPS dataset has an image size 16x16. The interpolation method used during resizing adjusts the pixel values to fit the larger image grid, enabling the model to be tested with standard image input sizes, such as 32x32. USPS image conversion is shown in Figure 6.
@@ 5.3 Experimental Setup
In this section, we describe the hyper-parameters and settings used for training the models. We are implementing our algorithm on the Pennylane quantum simulator. Hyper-parameters to run the algorithm were carefully selected to ensure stable training while optimizing performance for all the models.
@@ 5.3.1 Hyper-parameters
In our training process, several key hyper-parameters were used to ensure efficient learning and model convergence. The learning rate was set to $0.001$ for all the models. This learning rate controls how much the model’s weights are updated in response to the computed gradients, providing a stable training process while allowing sufficient progress over the course of the training. The models were trained for 200 epochs, with each epoch representing one full pass of the data set through the model. This duration was chosen to ensure that the models had enough time to learn the underlying features of the data, while regular monitoring of the validation loss helped prevent over-fitting. A batch size of $400$ was used, which means that $400$ images were processed together before updating the model parameters. This batch size was selected to balance computational efficiency and the stability of gradient estimation, helping smooth gradient updates and ensuring more stable training. For optimization, we used the Adam optimizer, which is known for its adaptive learning rate that adjusts based on the first and second moments of the gradients. The Adam optimizer is particularly effective for models with many parameters and large datasets. The loss function used was the Binary Cross-Entropy Loss (BCELoss), with the argument reduction = ’sum’, which ensures that the individual losses across all pixels are summed rather than averaged. This choice emphasizes the total error in the reconstruction of the images. To visualize the progress of the model over time, the variables were used to store a grid of reconstructed images at various epochs. These images allowed us to visually assess how well the models were reconstructing the images as training progressed. These configurations were chosen to ensure effective model training, allowing the models to learn from the data efficiently and perform well in both reconstruction and evaluation tasks.
6 FID and MSE Performance Evaluation and Discussion
To evaluate the performance of the proposed models, we used a combination of several key metrics: the Fréchet Inception Distance (FID) score[18], training and testing loss, and Mean Squared Error (MSE)[19].
Each of these metrics provides distinct insights into the model’s ability to generate better-quality images and generalize to new data: Fréchet Inception Distance (FID) Score measures the similarity between the distribution of real and generated images. A lower FID score indicates that the generated images are closer in distribution to the real images, reflecting higher image fidelity and quality. The FID score is particularly useful for assessing the realism of generated images in generative models. Mean Squared Error (MSE) measures the average squared difference between the generated and real images. A lower MSE indicates that the generated images are closer to the real images in terms of pixel-wise accuracy. This metric is commonly used to evaluate the reconstruction quality in generative models, with lower values indicating better performance. By combining these metrics, we gain a comprehensive understanding of how well our models perform in terms of both image quality and generalizability.
The FID score was computed for 1) images reconstructed from the input images and 2) images generated from random noise, giving us insight into how well the models perform in generating realistic images from these two different sources of input.
The results are shown in table 1 below:
@@ 6 FID and MSE Performance Evaluation and Discussion
Q-USPS 38.5 57.6 0.0058
The performance of all the models on the MNIST dataset was assessed by calculating the FID score using latent space representations as shown in Figure 7:
@@
For MNIST, the C-VAE achieves an FID score of 40.7 , indicating a reasonable ability to reconstruct images. CDP-VAE, demonstrates a slight improvement over the C-VAE with an FID score of 39.7 and then to Q-VAE, further improving upon CDP-VAE. The Q-VAE leverages quantum gates such as RY rotations to encode the data into a quantum state and Pauli-Z measurements bring quantum state to classical data, achieving a superior FID score of 37.3. This improved score suggests that the quantum model produces more faithful image reconstructions, capturing complex relationships in the data more effectively than the classical model. Although CDP-VAE approach is less complex, it still lacks the advanced feature extraction capabilities of the Q-VAE, which ultimately leads to lower reconstruction fidelity.
For the USPS dataset, the Q-VAE achieved an impressive FID score of 57.6, indicating better-quality image reconstructions and strong performance in terms of both fidelity and distribution similarity to the original dataset. This result highlights the effectiveness of the Q-VAE in handling simpler grayscale datasets, where it demonstrates superior feature extraction and reconstruction capabilities.
@@
Despite its simplicity, the CDP-VAE still achieved relatively good performance on the USPS dataset, with an FID score of 66.1 as shown in Figure 8.
On the MNIST dataset, C-VAE, with an MSE of 0.0093, showed slightly higher error compared to the Q-VAE, indicating that while it performs well, its feature extraction capabilities are not as advanced as those of the quantum model. Despite its relatively low MSE, the classical model’s performance is constrained by the limitations of traditional neural network-based encoding and decoding methods, which cannot capture the full complexity of the data as efficiently as quantum circuits. The CDP-VAE, with an MSE of 0.0092, performs little better then C-VAE While Q-VAE achieved the lowest MSE of 0.0088, demonstrating its ability to produce high-fidelity reconstructions compared to the other models. This is consistent with the Q-VAE’s superior performance across other metrics such as training and testing loss. The quantum-based encoding method enables the Q-VAE to extract richer feature representations, leading to more accurate reconstructions. The quantum model’s ability to capture complex dependencies between pixels allows it to outperform classical approaches, even when dealing with highly structured images like those in the MNIST dataset.